# A Diffusion-based Xray2MRI Model: Generating Pseudo-MRI Volumes From one Single X-ray

### v2

@@ 3.1 Employed database
3.2 Data preprocessing
3.2.1 X-ray
As shown in Fig. 3, the knee joint (Fig. 3) was identified from the plain radiograph (Fig. 3) through the YOLOv2 learning model [29], with images sized of 299 $\times$ 299 $\times$ 1, which were resized to the size of 256$\times$256$\times$1, and image intensity distribution was normalized to [-1, 1].
Figure 3: A standard knee plain radiograph and an identified knee joint highlighted in the red box 3. An identified knee joint 3.
3.2.2 MRI
@@
Figure 4: Distribution of the number of slices for each type of MRI sequence 4. Grey value distribution at each proportional position of slices within the T1-weighted MRI sequences 4.
After the processing, the slice count was fixed at 50 for the T1-weighted MRI scans as the final experimental dataset. The resulting 4,262 X-ray-MRI pairs were then randomly divided into training and validation sets with a ratio of 7:3.
3.3 Experimental details
The diffusion module utilized pre-trained weights from Zero-123 [30]. The pre-trained KOA classifier $\mathcal{C}$ and encoder $\mathcal{E}_{2}$ were sourced from the [31] and [32], respectively. During the training, a base learning rate of 1e-06 was applied, with a lambda linear scheduler incorporating a warm-up period every 100 steps. The batch size was set to 64, with $T$ configured to 1,000 steps, and an Exponential Moving Average (EMA) [33] of 0.99 was used for the diffusion sampling process. The scale factor of the latent space was set as 0.2. The AdamW optimizer [34] was utilized for training over 5,000 epochs. We implemented our approach using PyTorch v1.8.1 [35] on Nvidia A100 graphics cards with 80 GB of memory.
4 Results and discussion
4.1 Comparisons with State-of-the-Art Methods
In this section, the comparison is divided into two main parts: Qualitative visualization analysis and Quantitative metric-based analysis.
4.1.1 Qualitative visualization analysis
Table 1 provides a comprehensive visual comparison of corresponding pseudo-MRI slices generated by various methods, using X-ray images as input. To facilitate a clear and focused analysis, we selected three representative stages of KOA: G-0, indicating healthy knees; G-2, reflecting early-stage KOA; and G-4, corresponding to late-stage KOA. The compared methods include classic LDM [36], Cola-diff [18], X-diff [37], and our proposed approach. Each row displays the X-ray input on the far left, followed by the real MRI slice and the pseudo-MRI slices generated by the respective methods. To provide a more intuitive illustration of each method’s performance, difference maps generated using a Sobel operator filter are shown beneath each slice, highlighting pixel-wise deviations from the ground truth MRI by detecting edge differences. The visual results consistently show that our model produces pseudo-MRI slices that are closer to the ground truth, with significantly fewer and less pronounced deviations compared to other methods. Specifically, LDM suffers from the lack of constraints provided by the input X-ray. Without this conditional guidance, the model struggles to capture the detailed structural information of the knee joint, leading to great deviations from the ground truth MRI. On the other hand, compared with LDM, although Cola-Diff and X-Diff generate bone structures that are closer to the real MRI in terms of overall shape, they both lack a significant amount of fine details. This limitation prevents them from accurately capturing and reproducing the KOA symptoms. Moreover, it is noteworthy that it is particularly evident as the KOA grade of the input X-ray increases. The progressive narrowing of the knee joint space and the formation of osteophytes are learned and generated more accurately by our model, which demonstrates that our method not only ensures the generation of high-quality pseudo-MRI images but also maintains clinical relevance. However, the model’s predictions in areas outside the skeletal regions are less accurate due to the absence of critical information in X-ray images such as muscle and fat distribution, which will be discussed in Section 4.3.3.
Table 1: Visualization of a representative generated pseudo-MRI slice
@@ 4.1.1 Qualitative visualization analysis
    Here, to make it more intuitive and convenient, we only display a single slice from each MRI sequence that best represents the complete knee joint.
4.1.2 Quantitative metric-based analysis
@@
Figure 5: The box plots visualize the different performance of the evaluated approaches using PSNR 5 and SSIM/RSSIM 5.
4.2 Impact of the number of inference steps
@@
For the inference steps $s$, in addition to adhering to the original depth setting (i.e., $s=S=50$), we also experimented with varying inference steps (i.e., 30, 100, 200, 300, 400, and 500), to explore the impact of increasing the number of inference steps on the quality and accuracy of the generated slices. Fig. 6 shows the adjacent slice correlation coefficients under different inference steps. As can be seen, the red horizontal lines at 0.73 and 0.82 represent the mean adjacent slice correlation coefficients of the pseudo-MRI matching the volume of the original MRI (i.e., $s=S=50$) and that of the original MRI itself, respectively, with the latter serving as the upper benchmark for comparison. At 30 inference steps, the adjacent slice correlation starts at approximately 0.66, reflecting the impact of downsampling, which introduces inconsistencies and reduces the alignment between neighbouring slices. As the number of inference steps increases to 50, where the depth aligns with that of the original MRI volume, the correlation coefficient jumps sharply to around 0.73, suggesting a significant improvement in slice coherence. At 100 steps, the correlation continues to increase modestly. Beyond 100 steps, the improvement becomes more gradual as the number of inference steps increases, with diminishing returns, eventually stabilizing around 0.76 by 500 inference steps.
@@
On the other hand, for the sake of clarity and conciseness, we limit our visualization to the interpolation at $s=500$. Table 2 illustrates the synthesized intermediate frames between the original adjacent images (i.e., starting and ending images). Two distinct representative stages were evaluated: The transition from predominantly muscle tissue to the initial appearance of bone, and the shift from cartilage to the developing coronal knee joint. In each of these stages, the synthesized intermediate frames display relatively smooth and continuous transitions between bone and soft tissue structures. Importantly, the model successfully maintains key anatomical details throughout the transition, including the integrity of the bones, cartilage, and joint spaces, which underscores the robustness and effectiveness of these interpolation frames.
Table 2: Visualization of the synthesized intermediate frames
Starting image (real) $\stackrel{{\scriptstyle\text{Synthesized intermediate frames}}}{{\makebox[284.52756pt]{\rightarrowfill}}}$ Ending image (real)
The above results demonstrate that increasing the number of inference steps $s$ enhances interpolation, showcasing its potential for improving temporal resolution. By generating additional intermediate frames from a smaller set of real frames, interpolation effectively offers more continuous data without lengthening the actual scan duration. On the other hand, despite the improvements observed, the correlation coefficient from the interpolated pseudo-MRI fails to reach the 0.82 benchmark of the original MRI volume, even with the maximum 500 inference steps, which indicates that, while the interpolation process can enhance the consistency between adjacent slices, inherent limitations remain. These limitations arise from the inherent constraints of using X-rays as the sole source of information, which will be discussed in Section 4.3.3.
4.3 Discussion
In this paper, we introduce a novel diffusion-based Xray2MRI model that generates full pseudo-MRI volumes from a single X-ray image, representing one promising initial step toward cost-effective medical imaging solutions. The model leverages conditional inputs from X-rays, combined with target depth, KOA probability distribution, and image intensity distribution modules, to guide the synthesis process and ensure anatomically accurate MRI slice generation. Experimental results show that the generated pseudo-MRI sequences approximate real MRI scans, delivering high image quality while preserving critical clinical relevance. Additionally, by increasing the number of inference steps, the model exhibits effective interpolation, resulting in smoother transitions between slices. Several points are to be discussed.
4.3.1 How it works?
@@
4.3.2 Why use a 2D diffusion model instead of a 3D one directly?
@@
We opted to use a 2D diffusion model, generating individual MRI slices at different target depths during the inference phase, which are then stacked to form a pseudo-3D MRI sequence. By using this method, the model can effectively learn and capture the spatial relationships between the individual slices, allowing it to understand inter-slice dependencies. Additionally, employing a 2D model reduces computational complexity, making both the training and inference processes more efficient compared to handling full 3D data. This approach also enhances the interpretability of the entire process, as it enables a more granular analysis of how each slice contributes to the final 3D reconstruction.
4.3.3 Strengths and limitations
@@
This study has several notable strengths. Firstly, it introduces an innovative method for generating pseudo-MRI images from one single X-ray image, attempting to bridge the gap between these two imaging modalities using advanced AI technology. This approach represents a novel attempt in cases where MRI may not be readily available. Our experiments demonstrate that incorporating supplementary information beyond the X-ray images themselves, such as disease probability distributions, enhances the quality of the generated pseudo-MRI, which underscores the potential of integrating more diverse information sources in the future to produce more accurate pseudo-MRI volumes. Second, the study employs a 2D diffusion model, which not only reduces the computational resources required but also allows for more detailed control over the generated images, leading to a more interpretable and transparent reconstruction process. Finally, the study’s focus on clinically relevant metrics, including both traditional image quality assessments and pathology-specific evaluations, ensures that the generated pseudo-MRI images are not only visually similar to ground truth MRIs but also clinically relevant. This study has several limitations that warrant consideration. First, our model are highly dependent on the availability of large, high-quality paired datasets. The scarcity of such datasets, particularly in certain clinical scenarios or for rare diseases, may limit the generalizability of our approach. Second, although our model is able to approximate MRI features from X-ray inputs, it remains strongly uncertain whether these pseudo-MRI images can be reliably used for clinical diagnosis without the support of additional imaging modalities as the model cannot learn information that is not present in the input X-ray image, which is precisely the value that MRI imaging provides. Such limitation is clearly illustrated in Fig. 7. Finally, the computational resources required for training, despite being reduced by using a 2D model, high-performance hardware is still necessary, which presents a challenge for broader implementation in clinical settings.
(a) X-ray
(b) Real MRI
(c) Generated MRI
Figure 7: The X-ray 7(a) does not provide information about cartilage abnormalities, the generated pseudo-MRI 7(c) thus similarly lacks this detail, whereas such information is presented in the real MRI 7(b).
4.3.4 Future work
More comprehensive information (e.g., BMI, age, gender, disease type and location, etc) could be effectively integrated with X-ray images to further enhance the generation performance. Moreover, to improve the model’s deployability, the use of knowledge distillation could be of interest.
5 Statements
@@
6 Acknowledgements
The authors gratefully acknowledge the support of the Ralph Schlaeger Research Fellowship under award number 2024A018081 from Massachusetts General Hospital (MGH), Harvard Medical School (HMS).

### v3

@@ III-A Employed database
III-B Data preprocessing
III-B1 X-ray
As shown in Fig. 3, the knee joint (Fig. 3b) was identified from the plain radiograph (Fig. 3a) through the YOLOv2 learning model [28], with images sized of 299 $\times$ 299 $\times$ 1, which were resized to the size of 256$\times$256$\times$1, and image intensity distribution was normalized to [-1, 1].
Figure 3: A standard knee plain radiograph and an identified knee joint highlighted in the red box 3a. An identified knee joint 3b.
III-B2 MRI
@@
Figure 4: Distribution of the number of slices for each type of MRI sequence 4a. Grey value distribution at each proportional position of slices within the T1-weighted MRI sequences 4b.
III-B3 Radiographic features
The encoding process for the radiographic features involves transforming each feature into a One-Hot vector representation based on its grades. For missing values, all positions in the One-Hot vector are set to zero to indicate the absence of valid information for that feature, which ensures that the model can distinguish missing values from valid inputs without introducing artificial bias. The resulting One-Hot vectors for all radiographic features are concatenated to form a unified representation, which is subsequently used as part of the guidance input during the denoising process in the diffusion model.
After the above processing, the slice count was fixed at 50 for the T1-weighted MRI scans as the final experimental dataset. The resulting 4,262 X-ray-MRI pairs with respective radiographic feature information were then randomly divided into training, validation and testing sets with a ratio of 7:2:1.
III-C Experimental details
The diffusion module utilized pre-trained weights from Zero-123 [29]. The pre-trained encoder $\mathcal{E}_{2}$ is sourced from [30]. During the training, a base learning rate of 1e-06 was applied, with a lambda linear scheduler incorporating a warm-up period every 100 steps. The batch size was set to 64, with $T$ configured to 2,000 steps, and an Exponential Moving Average (EMA) [31] of 0.99 was used for the diffusion sampling process. The scale factor of the latent space was set as 0.2. The AdamW optimizer [32] was utilized for training over 5,000 epochs. We implemented our approach using PyTorch v1.8.1 [33] on Nvidia A100 graphics cards with 80 GB of memory.
IV Results and discussion
TABLE II: Visualization of a representative generated MRI slice
@@ IV Results and discussion
    Here, to make it more intuitive and convenient, we only display one slice from each MRI sequence that best represents the complete knee joint.
IV-A Comparisons with State-of-the-Art Methods
The comparison is divided into two main parts: Qualitative visualization analysis and quantitative metric-based analysis.
IV-A1 Qualitative visualization analysis
Table II provides a comprehensive visual comparison of corresponding generated MRI slices generated by various methods, using X-ray images as input. To facilitate a clear and focused analysis, we selected three representative stages of KOA: G-0, indicating healthy knees; G-2, reflecting early-stage KOA; and G-4, corresponding to late-stage KOA. The compared methods include classic LDM [34], Cola-diff [18], X-diff [35], and our proposed approach. Each row displays the X-ray input on the far left, followed by the real MRI slice and the MRI slices generated by the respective methods. To provide a more intuitive illustration of each method’s performance, difference maps generated using a Sobel operator filter are shown beneath each slice, highlighting pixel-wise deviations from the ground truth MRI by detecting edge differences. The visual results consistently show that our model produces generated MRI slices that are closer to the ground truth, with significantly fewer and less pronounced deviations compared to other methods. Specifically, LDM suffers from the lack of constraints provided by the input X-ray. Without this conditional guidance, the model struggles to capture the detailed structural information of the knee joint, leading to great deviations from the ground truth MRI. On the other hand, compared with LDM, although Cola-Diff and X-Diff generate bone structures that are closer to the real MRI in terms of overall shape, they both lack a significant amount of fine details. This limitation prevents them from accurately capturing and reproducing the KOA symptoms. Moreover, it is noteworthy that it is particularly evident as the KOA grade of the input X-ray increases. The progressive narrowing of the knee joint space and the formation of osteophytes are learned and generated more accurately by our model, which demonstrates that our method not only ensures the generation of high-quality MRI but also maintains clinical relevance. However, the model’s predictions in areas outside the skeletal regions are less accurate due to the absence of critical information in X-ray images such as muscle and fat distribution, which will be discussed in Section IV-D3.
IV-A2 Quantitative metric-based analysis
@@
Figure 5: The box plots visualize the different performance of the evaluated approaches using PSNR 5a and SSIM/RSSIM 5b.
IV-B Ablation study on the feature information
To assess the contribution of the integrated patient-specific radiographic feature information $r$ within the guidance module, we conducted a detailed ablation study, analyzing the impact of specific variables within $r$. This ablation study was performed using the same dataset configuration to ensure consistency and comparability across experiments. Here, we particularly focused on the roles of the variables of chondrocalcinosis (CPPD), attrition, and cysts in Table I, given their significant clinical relevance and the challenges associated with interpreting these features from X-ray images alone. For instance, CPPD is associated with the deposition of calcium pyrophosphate crystals in the joint tissues, which manifests as subtle calcifications, which often overlap with normal joint structures, leading to ambiguity in detection using Xrays. Attrition, on the other hand, reflects the loss of articular cartilage and underlying bone, a hallmark of joint degeneration. Its appearance on X-rays can be subtle. Similarly, cystic changes represent fluid-filled cavities or bone erosions that occur in advanced joint diseases, these changes can appear as radiolucent areas on X-rays, and their true extent and clinical implications are difficult to discern without complementary information. By isolating these variables, we aimed to determine the extent to which they enhance the model’s ability to synthesize MRI outputs that accurately reflect the underlying pathology.
As shown in Table III, before integrating the CPPD variable (V00XRCHL), the model generated MRI sequences that entirely lacked any representation of CPPD, as it was unable to incorporate this feature without explicit guidance. However, after integrating CPPD as a guidance variable, the model began to display CPPD in a roughly accurate region (i.e., the lateral compartment). For the attrition, integrating the corresponding variable (V00XRATTL) enabled the model to better capture signs of cartilage loss and bone alterations in the tibia lateral compartment. Similarly, for the cysts, the variable (V00XRCYTL) allowed the model to successfully synthesize fluid-filled regions in the tibia lateral compartment. It is noteworthy that, while the model struggled to fully capture the precise severity and spatial details of certain features, the ablation studies demonstrate that integrating patient-specific radiographic information beyond what X-rays alone provide improves the anatomical realism and clinical relevance of the generated MRI sequences, allows the model to generate MRI outputs that are, to some extent, closer to real MRI scans, which underscores the immense potential of incorporating additional information (e.g., Body Mass Index (BMI), neural distribution, etc) as guidance into such models.
TABLE III: Ablation studies for feature information
Ground Truth with CPPD1 without CPPD
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth with Cysts3 without Cysts
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth with Attrition2 without Attrition
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
  • $1$
    V00XRCHL=1: CPPD present in lateral compartment
  • $2$
    V00XRCYTL=1: Cysts present in the tibia lateral compartment
  • $3$
    V00XRATTL=1: Attrition present in tibia lateral compartment
TABLE IV: Visualization of the synthesized intermediate frames
Starting image (real) $\stackrel{{\scriptstyle\text{\footnotesize Synthesized intermediate frames}}}{{\makebox[284.52756pt]{\rightarrowfill}}}$ Ending image (real)
IV-C Impact of the number of inference steps
@@
For the inference steps $s$, in addition to adhering to the original depth setting (i.e., $s=S=50$), we also experimented with varying inference steps (i.e., 30, 100, 200, 300, 400, and 500), to explore the impact of increasing the number of inference steps on the quality and accuracy of the generated slices. Fig. 6 shows the adjacent slice correlation coefficients under different inference steps. As can be seen, the red horizontal lines at 0.73 and 0.82 represent the mean adjacent slice correlation coefficients of the generated MRI matching the volume of the original MRI (i.e., $s=S=50$) and that of the original MRI itself, respectively, with the latter serving as the upper benchmark for comparison. At 30 inference steps, the adjacent slice correlation starts at approximately 0.66, reflecting the impact of downsampling, which introduces inconsistencies and reduces the alignment between neighbouring slices. As the number of inference steps increases to 50, where the depth aligns with that of the original MRI volume, the correlation coefficient jumps sharply to around 0.73, suggesting a significant improvement in slice coherence. At 100 steps, the correlation continues to increase modestly. Beyond 100 steps, the improvement becomes more gradual as the number of inference steps increases, with diminishing returns, eventually stabilizing around 0.76 by 500 inference steps. Table IV illustrates the synthesized intermediate frames between the original adjacent images (i.e., starting and ending images) using 500 inference steps. A distinct representative stage is evaluated: The transition from predominantly muscle tissue to the initial appearance of bone. In this stage, the synthesized intermediate frames display relatively smooth and continuous transitions between bone and soft tissue structures, which underscores the robustness and effectiveness of these interpolation frames.
@@
The above results demonstrate that increasing the number of inference steps $s$ enhances interpolation, showcasing its potential for improving temporal resolution. By generating additional intermediate frames from a smaller set of real frames, interpolation effectively offers more continuous data without lengthening the actual scan duration. On the other hand, despite the improvements observed, the correlation coefficient from the interpolated MRI fails to reach the 0.82 benchmark of the original MRI volume, even with the maximum 500 inference steps, which indicates that, while the interpolation process can enhance the consistency between adjacent slices, inherent limitations remain. These limitations arise from the inherent constraints of using X-rays as the sole imaging source of information, which will be discussed in Section IV-D3.
IV-D Discussion
In this paper, we introduce a novel diffusion-based model that generates full MRI volumes from a single X-ray image. Specifically, the model leverages conditional inputs from X-rays, combined with target depth and patient-specific feature information, to guide the synthesis process for the MRI slice generation. Experimental results demonstrate that, compared to SOTA methods, our approach generates pseudo-MRI sequences that are closer to real MRI scans in terms of image quality and clinical relevance. Furthermore, ablation studies on the guidance module validate that incorporating supplementary information beyond what is available from X-rays alone improves the accuracy of the generated MRI, which highlights the immense potential for integrating additional independent information in the future to further enhance MRI generation performance. Lastly, by increasing the number of inference steps, the model exhibits effective interpolation, resulting in smoother transitions between slices. Several points are to be discussed.
IV-D1 How it works?
@@
IV-D2 Why use a 2D diffusion model instead of a 3D one directly?
@@
We opted to use a 2D diffusion model, generating individual MRI slices at different target depths during the inference phase, which are then stacked to form a 3D MRI sequence. By using this method, the model can effectively learn and capture the spatial relationships between the individual slices, allowing it to understand inter-slice dependencies. Additionally, employing a 2D model reduces computational complexity, making both the training and inference processes more efficient compared to handling full 3D data. This approach also enhances the interpretability of the entire process, as it enables a more granular analysis of how each slice contributes to the final 3D reconstruction.
IV-D3 Strengths and limitations
@@
This study has several notable strengths. Firstly, it introduces an innovative method for generating MRI from one single X-ray image, attempting to bridge the gap between these two imaging modalities using advanced AI technology, which represents a novel attempt in cases where MRI may not be readily available. Our experiments demonstrate that incorporating supplementary information (i.e., patient-specific radiographic feature information in this study) beyond the X-ray images themselves enhances the quality of the generated MRI, which underscores the potential of integrating more diverse information sources in the future to produce more accurate generated MRI volumes. Second, the study employs a 2D diffusion model, which not only reduces the computational resources required but also allows for more detailed depth control over the generated images, leading to a more interpretable and transparent reconstruction process. Finally, the study’s focus on clinically relevant metrics, including both traditional image quality assessments and pathology-specific evaluations, ensures that the generated MRI are not only visually similar to ground truth but also clinically relevant. This study has several limitations that warrant consideration. Our model are highly dependent on the availability of large, high-quality paired datasets. The scarcity of such datasets, particularly in certain clinical scenarios or for rare diseases, may limit the generalizability of our approach. Moreover, the computational resources required for training, despite being reduced by using a 2D model, high-performance hardware is still necessary, which presents a challenge for broader implementation in clinical settings. Therefore, more additional information (e.g., Body Mass Index (BMI), age, gender, etc) could be integrated to further enhance the generation performance. To improve the model’s deployability, the use of knowledge distillation could be of interest.
V Statements
@@
VI Acknowledgements
The authors gratefully acknowledge the support of the Ralph Schlaeger Research Fellowship under award number 246448 from Massachusetts General Hospital (MGH), Harvard Medical School (HMS) and the French National Agency of Research (ANR) under project number ANR-20-CE45-0013-01.